# Increasing Batch Size Improves Convergence of Stochastic Gradient Descent with Momentum

### v1

@@ A.1 Proposition and Lemma
$=\left\|\frac{1-\beta}{1-\beta^{t}}\beta\left(\beta^{t-1}\nabla f(\bm{\theta}_{0})+\beta^{t-2}\nabla f(\bm{\theta}_{1})+\cdots+\beta^{t-(t-1)}\nabla f(\bm{\theta}_{t-2})+\nabla f(\bm{\theta}_{t-1})\right)-\frac{1-\beta^{t-1}}{1-\beta^{t}}\beta\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
@@
$\|\boldsymbol{m}_{t-1}\rVert^{2}$ $\leq 2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\Bigg{\{}2(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
@@
$=2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+4(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
@@
$-\eta X_{t}$ $\leq\frac{\rho\eta^{3}L^{2}}{2}\left(\frac{\beta}{1-\beta}\right)^{2}\Bigg{\{}2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+4(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
@@
$=\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}(1-\beta^{t})^{2}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$
@@
$\begin{split}-\eta\mathbb{E}[X_{t}]&\leq\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta^{t-1})^{2}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}(1-\beta^{t})^{2}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right]+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\leq\rho\eta^{3}L^{2}\beta^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}{(1-\beta^{t})^{2}}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right]+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}].\end{split}$ (12)
@@
Therefore, from (7), (12), and (13), for all $t\geq 2$ and for all $\rho>0$,
$\mathbb{E}[f(\boldsymbol{z}_{t+1})]$ $\leq\mathbb{E}[f(\boldsymbol{z}_{t})]-\eta\mathbb{E}[X_{t}]+\frac{L\eta^{2}}{2}\mathbb{E}[Y_{t}]$
$\leq\mathbb{E}[f(\boldsymbol{z}_{t})]+\rho\eta^{3}L^{2}\beta^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}{(1-\beta^{t})^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}+\frac{L\eta^{2}}{2}\left(\frac{\sigma^{2}}{b_{t}}+\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\right)$
$=\mathbb{E}[f(\boldsymbol{z}_{t})]+\underbrace{\left\{2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}+\eta\left(\frac{1}{2\rho}-1\right)+\frac{L\eta^{2}}{2}\right\}}_{A}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
@@
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
$L_{t}:=\begin{cases}f(\boldsymbol{z}_{0})-f^{\star}&\text{ }(t=0)\\ f(\boldsymbol{z}_{1})-f^{\star}&\text{ }(t=1)\\ f(\boldsymbol{z}_{t})-f^{\star}+\sum_{i=1}^{t-1}c_{i}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}&\text{ }(t\geq 2),\end{cases}$
@@
$\mathbb{E}[L_{t+1}-L_{t}]$ $=\mathbb{E}[f(\boldsymbol{z}_{t+1})-f(\boldsymbol{z}_{t})]+\mathbb{E}\left[\sum_{i=1}^{t}c_{i}\left\|\boldsymbol{\theta}_{t+2-i}-\boldsymbol{\theta}_{t+1-i}\right\rVert^{2}-\sum_{i=1}^{t-1}c_{i}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}\right]$
$=\mathbb{E}[f(\boldsymbol{z}_{t+1})-f(\boldsymbol{z}_{t})]+\mathbb{E}[c_{1}\left\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}_{t}\right\rVert^{2}]+\mathbb{E}\left[\sum_{i=1}^{t-1}c_{i+1}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}-\sum_{i=1}^{t-1}c_{i}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}\right]$
$=\mathbb{E}[f(\boldsymbol{z}_{t+1})-f(\boldsymbol{z}_{t})]+\mathbb{E}[c_{1}\left\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}_{t}\right\rVert^{2}]+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}].$
@@
$\quad+c_{1}\eta^{2}\mathbb{E}[\left\|\boldsymbol{m}_{t}\right\rVert^{2}]+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}].$ (14)
A similar discussion to the one for showing (9) and (10) ensures that
$\|\boldsymbol{m}_{t}\rVert^{2}$ $\leq 2\left\|\boldsymbol{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}$
@@
$\mathbb{E}[\|\bm{m}_{t}\|^{2}]$ $\leq 2(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}+4(1-\beta^{t})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}+4(1-\beta^{t})^{2}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$\leq 2(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}+4\|\nabla f(\bm{\theta}_{t})\|^{2}+4(1-\beta^{t})^{2}\underbrace{\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}}_{V_{t}}.$
Lemma 2 in (Liu et al., 2020) guarantees that
@@
$\mathbb{E}[L_{t+1}-L_{t}]$ $\leq\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
@@
$\quad+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}]$
@@
$\quad+\sum_{i=1}^{t-1}\underbrace{\left\{\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}a_{t,t-i}+4c_{1}\eta^{2}(1-\beta^{t})^{2}a_{t,t-i}+(c_{i+1}-c_{i})\right\}}_{N_{t,i}}\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}].$
@@
=$=\left\{\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}+4c_{1}\eta^{2}(1-\beta^{t})^{2}\right\}\frac{L^{2}\beta^{i}}{1-\beta^{t}}\left(i+\frac{\beta}{1-\beta}\right)-\left(4c_{1}\eta^{2}+\frac{L\eta^{2}}{{(1-\beta)^{2}}}\right)\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)L^{2}$
@@
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
where $\beta\in[0,1)$, and $\eta$ and $c_{1}$ satisfy (16). Summing the above inequality from $t=0$ to $t=T-1$ gives that
@@
Moreover, from (18), we have that
@@
where $T=\sum_{m=0}^{M}K_{m}E_{m}$. Define $K_{\max}:=\max\{K_{m}\colon m\in[0:M]\}$ and $E_{\max}:=\max\{E_{m}\colon m\in[0:M]\}$. Then, we have that
@@
Proof of Theorem 3.4.
NSHB with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2) (Section 2.2). Hence, Theorem 3.3 leads to Theorem 3.4. ∎
A.3 Proofs of Theorems 3.1 and 3.2
Using Lemma A.2 and the proof of Lemma A.3 with $\rho=\frac{1-\beta}{2L\eta}$, we have the following lemma. Hence, we omit the proof of Lemma A.6.
@@
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1. Let $(\bm{z}_{t})$ be the sequence defined as in Lemma A.3. Then, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}[f(\boldsymbol{z}_{t+1})]$ $\leq\mathbb{E}[f(\boldsymbol{z}_{t})]+\eta\left\{L\left(\frac{1+\beta^{2}}{1-\beta}+\frac{1}{2}\right)\eta-1\right\}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
@@
Using Lemma A.6 and the proof of Lemma A.4 with $\rho=\frac{1-\beta}{2L\eta}$, we have the following lemma. Hence, we omit the proof of Lemma A.7.
@@
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
$\mathbb{E}[L_{t+1}-L_{t}]$ $\leq\eta\left\{\frac{(3-\beta+\beta^{2})L\eta}{2(1-\beta)}-1+4c_{1}\eta\right\}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+\left\{\left(\frac{\beta^{2}}{1+\beta}+1\right)\frac{L}{2}+\frac{2c_{1}(1-\beta)}{1+\beta}\right\}\frac{\eta^{2}\sigma^{2}}{b}.$
@@
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
Proofs of Theorems 3.1 and 3.2.
Lemma A.8 leads to the assertion in Theorem 3.1. NSHB with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2) (Section 2.2). Hence, Theorem 3.1 leads to Theorem 3.2. ∎

### v2

@@ A.1 Proposition and lemma
$=\left\|\frac{1-\beta}{1-\beta^{t}}\beta\left(\beta^{t-1}\nabla f(\bm{\theta}_{0})+\cdots+\beta^{t-(t-1)}\nabla f(\bm{\theta}_{t-2})+\nabla f(\bm{\theta}_{t-1})\right)-\frac{1-\beta^{t-1}}{1-\beta^{t}}\beta\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
@@
$\|\bm{m}_{t-1}\rVert^{2}$ $\leq 2\left\|\bm{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\Bigg\{2(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
@@
$=2\left\|\bm{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+4(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
@@
$-\eta X_{t}$ $\leq\frac{\rho\eta^{3}L^{2}}{2}\left(\frac{\beta}{1-\beta}\right)^{2}\Bigg\{2\left\|\bm{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+4(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
@@
$=\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\left\|\bm{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}(1-\beta^{t})^{2}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$\quad+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$
@@
$\begin{split}-\eta\mathbb{E}[X_{t}]&\leq\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}\\ &\quad+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta^{t-1})^{2}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}(1-\beta^{t})^{2}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right]\\ &\quad+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\leq\rho\eta^{3}L^{2}\beta^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}{(1-\beta^{t})^{2}}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right]\\ &\quad+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}].\end{split}$ (12)
@@
Therefore, from (7), (12), and (13), for all $t\geq 2$ and for all $\rho>0$, we have
$\mathbb{E}[f(\bm{z}_{t+1})]$ $\leq\mathbb{E}[f(\bm{z}_{t})]-\eta\mathbb{E}[X_{t}]+\frac{L\eta^{2}}{2}\mathbb{E}[Y_{t}]$
$\leq\mathbb{E}[f(\bm{z}_{t})]+\rho\eta^{3}L^{2}\beta^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}{(1-\beta^{t})^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$\quad+\frac{L\eta^{2}}{2}\left(\frac{\sigma^{2}}{b_{t}}+\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\right)$
$=\mathbb{E}[f(\bm{z}_{t})]+\underbrace{\left\{2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}+\eta\left(\frac{1}{2\rho}-1\right)+\frac{L\eta^{2}}{2}\right\}}_{A}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
@@
Suppose that Assumption 1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
$L_{t}:=\begin{cases}f(\bm{z}_{0})-f^{\star}&\text{ }(t=0)\\ f(\bm{z}_{1})-f^{\star}&\text{ }(t=1)\\ f(\bm{z}_{t})-f^{\star}+\sum_{i=1}^{t-1}c_{i}\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}&\text{ }(t\geq 2),\end{cases}$
@@
$\mathbb{E}[L_{t+1}-L_{t}]$ $=\mathbb{E}[f(\bm{z}_{t+1})-f(\bm{z}_{t})]+\mathbb{E}\left[\sum_{i=1}^{t}c_{i}\left\|\bm{\theta}_{t+2-i}-\bm{\theta}_{t+1-i}\right\rVert^{2}-\sum_{i=1}^{t-1}c_{i}\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}\right]$
$=\mathbb{E}[f(\bm{z}_{t+1})-f(\bm{z}_{t})]+\mathbb{E}[c_{1}\left\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\right\rVert^{2}]$
$\quad+\mathbb{E}\left[\sum_{i=1}^{t-1}c_{i+1}\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}-\sum_{i=1}^{t-1}c_{i}\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}\right]$
$=\mathbb{E}[f(\bm{z}_{t+1})-f(\bm{z}_{t})]+\mathbb{E}[c_{1}\left\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\right\rVert^{2}]$
$\quad+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}].$
@@
$\quad+c_{1}\eta^{2}\mathbb{E}[\left\|\bm{m}_{t}\right\rVert^{2}]+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}].$ (14)
A similar discussion to the one showing (9) and (10) ensures that
$\|\bm{m}_{t}\rVert^{2}$ $\leq 2\left\|\bm{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}$
@@
$\mathbb{E}[\|\bm{m}_{t}\|^{2}]$ $\leq 2(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}+4(1-\beta^{t})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+4(1-\beta^{t})^{2}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$\leq 2(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}+4\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+4(1-\beta^{t})^{2}\underbrace{\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}}_{V_{t}}.$
Lemma 2 in [Liu et al., 2020] guarantees that
@@
$\mathbb{E}[L_{t+1}-L_{t}]$
$\leq\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
@@
$\quad+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}]$
@@
$\quad+\sum_{i=1}^{t-1}\underbrace{\left\{\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}a_{t,t-i}+4c_{1}\eta^{2}(1-\beta^{t})^{2}a_{t,t-i}+(c_{i+1}-c_{i})\right\}}_{N_{t,i}}\mathbb{E}[\left\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\right\rVert^{2}].$
@@
$=\left\{\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}+4c_{1}\eta^{2}(1-\beta^{t})^{2}\right\}\frac{L^{2}\beta^{i}}{1-\beta^{t}}\left(i+\frac{\beta}{1-\beta}\right)$
$\quad-\left(4c_{1}\eta^{2}+\frac{L\eta^{2}}{{(1-\beta)^{2}}}\right)\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)L^{2}$
@@
Suppose that Assumption 1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
where $\beta\in[0,1)$, and $\eta$ and $c_{1}$ satisfy (16). Summing the above inequality from $t=0$ to t=T1t=T-1 gives
@@
Moreover, from (18), we have
@@
where $T=\sum_{m=0}^{M}K_{m}E_{m}$. Define $K_{\max}:=\max\{K_{m}\colon m\in[0:M]\}$ and $E_{\max}:=\max\{E_{m}\colon m\in[0:M]\}$. Then, we have
@@
Proof of Theorem A.2.
NSHB with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2) (Section 2.2). Hence, Theorem 3.2 leads to Theorem A.2. ∎
A.3 Proofs of Theorems 3.1 and A.1
Lemma A.1 and the proof of Lemma A.2 with $\rho=\frac{1-\beta}{2L\eta}$ immediately yield the following lemma. Hence, we will omit its proof.
@@
Suppose that Assumption 1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1. Let $(\bm{z}_{t})$ be the sequence defined as in Lemma A.2. Then, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}[f(\bm{z}_{t+1})]$ $\leq\mathbb{E}[f(\bm{z}_{t})]+\eta\left\{L\left(\frac{1+\beta^{2}}{1-\beta}+\frac{1}{2}\right)\eta-1\right\}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
@@
Moreover, Lemma A.5 and the proof of Lemma A.3 with $\rho=\frac{1-\beta}{2L\eta}$ yield the following lemma, whose proof we will also omit.
@@
Suppose that Assumption 1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
$\mathbb{E}[L_{t+1}-L_{t}]$ $\leq\eta\left\{\frac{(3-\beta+\beta^{2})L\eta}{2(1-\beta)}-1+4c_{1}\eta\right\}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+\left\{\left(\frac{\beta^{2}}{1+\beta}+1\right)\frac{L}{2}+\frac{2c_{1}(1-\beta)}{1+\beta}\right\}\frac{\eta^{2}\sigma^{2}}{b}.$
@@
Suppose that Assumption 1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
@@
Proof of Theorems 3.1 and A.1.
Lemma A.7 leads to the assertion in Theorem 3.1. NSHB with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2) (Section 2.2). Hence, Theorem 3.1 leads to Theorem A.1. ∎
Theorem A.1 (Upper bound of $\min_{t}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|$ of mini-batch SHB with Constant BS).
Suppose that Assumption 1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by (2) with a momentum weight $\beta\in(0,1)$, a constant learning rate $\alpha>0$ such that
$\alpha\leq\frac{(1-\beta)^{2}}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},$
and Constant BS defined by (3). Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(1-\beta)(f(\bm{\theta}_{0})-f^{\star})}{\alpha T}+\frac{L\alpha\sigma^{2}}{(1-\beta)b}\left\{\frac{3\beta^{2}+\beta}{2(1+\beta)}+1\right\},$
that is,
Theorem A.2 (Convergence of mini-batch SHB with Exponential Growth BS).
Suppose that Assumption 1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by (2) with a momentum weight $\beta\in(0,1)$, a constant learning rate $\alpha>0$ such that
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(1-\beta)(f(\bm{\theta}_{0})-f^{\star})}{\alpha T}+\frac{2L\alpha\sigma^{2}K_{\max}E_{\max}\delta}{(1-\beta)(\beta^{2}\delta-1)b_{0}T}\left(\frac{\beta^{2}}{1-\beta^{2}}-\frac{1}{\delta-1}\right),$
that is,